WARNING: This web is not used anymore. Please use PDBService.MinuteS13March08 instead!
 

Minutes 13 March 08

  • Phone:
    • Gridka: Doris and Andreas
    • RAL: Carmine
    • NDGF: Olli
    • BNL: Carlos
    • CNAF: Barbara
    • SARA: Alexander
    • TRIUMF: Andrew

  • CERN:
    • 3D: Dirk, Maria, Eva
    • ATLAS: Gancho, Florbela

  • General announcements:
    • Next workshop: scheduled as part of the WLCG workshop, draft agenda sent around by Maria, please send your comments
      • Will attend: Gordon and Carmine, Alexander, Carlos, Olli, Barbara and Federico, Doris and Andreas, Florbela and Gancho
      • Please send us suggestions for the Agenda
    • 3d project from the development point of view is completed, so Dirk will propose to conclude this project. Status report will be prepared for the Management Board.
      • Dirk will prepare a draft on how this will continue (operational meetings, 3d as part of the databases services)
    • Licenses review: Licenses discussion, possible common acquisition as in 2006, Dirk has asked the sites to send their requests.
      • Licenses from 2006 are still valid, can be moved to new hw, but taking in count new rules for multi-core hw.
      • From CERN it will be not possible to get another similar arrangement during some time.
      • Feedback must be collected soon because the acquisition will be done in April.

  • Status Tier0 and Tier1 sites:
    • Tier0:
      • Patch for bug: imported table do not replicated
      • ORA-600 error when dropping propagation fixed in 11.2, backport requested
      • Split performed for PIC and CNAF to be prepared for their interventions.
    • Triumf: CPU January patch will be applied today
    • SARA: Nothing to report. Meeting with DQ2 people to be scheduled.
    • BNL:
      • CPU patch FTS database.
      • Views recompilation on 3d cluster next week.
      • One of the nodes was rebooted by itself but service was not affected, bug fixed in 10.2.0.4
      • Dawid is helping Carlos with the 32bits to 64bits migration tests
    • RAL: Next week January CPU patch on both databases
    • NDGF: Patch applied next week
    • CNAF: Preparation of the big intervention: downtime for 1 week, after CPU patch will be applied
    • GridKa: LHCb/LFC frontend installation finished. CPU January to be applied

  • Contact GOC db so the EGEE broadcast is sent automatically
  • Experiments status:
    • PVSS replication to be moved to production soon

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-04-21 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback